The expression of the magnetic field associated with an electromagnetic wave in
vacuum is given by
B = (100T) y sin (2πx 10⁸t + kz)
Determine the wave number, frequency and the direction of propagation of the wave
and the magnitude and direction of the electric field associated with it.
1
Expert's answer
2020-05-12T09:57:41-0400

As per the given question,

"B=(100T)y\sin(2π×10^8t+kz)"

Wave number is given = K


frequency ="10^8Hz"

Direction of the propagation is in along the z axis

We know that the relation between the electric field (E)and the magnetic field(B) and the speed of light(c) is

"c=\frac{E}{B}"


"E=cB=3\u00d710^8\u00d7100y=3\u00d710^{10}N\/c"